# Crossed products by compact group actions with the weak tracial Rokhlin property

### v2

@@ Abstract.
In this paper, we introduce compact group actions with the weak tracial Rokhlin property. This concept generalizes both the finite group actions with the weak tracial Rokhlin property and the compact group actions with the tracial Rokhlin property on classifiable C*-algebras (in the sense of the Elliott program). Under this framework, we prove that simplicity, pure infiniteness, tracial $\mathcal{Z}$-stability and the combination of nuclearity and $\mathcal{Z}$-stability can be transferred from the original algebra to the crossed product. We also show that the radius of comparison of the fixed point algebra does not exceed that of the original algebra. Furthermore, we discuss the relationship between our definitions and natural generalizations of the finite group case in non-Elliott program settings. Finally, we provide a nontrivial example of a compact group action with the weak tracial Rokhlin property with comparison: an action of $(S_{2})^{\mathbb{N}}$ on the Jiang-Su algebra $\mathcal{Z}$. Since $\mathcal{Z}$ contains no nontrivial projections, this action does not possess the tracial Rokhlin property.
@@ 1. Introduction
The Rokhlin property for the case of a single automorphism was originally introduced for von Neumann algebras by Connes in [9]. Later, the Rokhlin property for finite group actions on C*-algebras first appeared in the work of Herman and Jones in [26] and [27]. This property is useful to understand the structure of the crossed product of C*-algebras and properties passing from the original algebra to the crossed product [39]. However, the finite group acitions with the Rokhlin property are rare. Phillips, in [41], introduced the tracial Rokhlin property for finite group actions on unital simple C*-algebras. The tracial Rokhlin property is generic in many cases, and also can be used to study properties passing from the original algebra to the crossed product. Weak versions of the tracial Rokhlin property in which one uses orthogonal positive contractions instead of orthogonal projections were studied for actions on unital simple C*-algebras with few projections [20, 28, 36, 42, 52, 53] (see Definition 3.1). As an example, the flip action on the Jiang-Su algebra $\mathcal{Z}\cong\mathcal{Z}\otimes\mathcal{Z}$ has the weak tracial Rokhlin property but it does not have the tracial Rokhlin property [28]. For the non-unital case, Santiago and Gardella studyed the Rokhlin property for finite group actions on non-unital simple C*-algebras in [47] and [22]. Forough and Golestani studyed the (weak) tracial Rokhlin property for finite group actions on non-unital simple C*-algebras in [13].
@@ 2.1. Central sequence algebras
Let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of $G$ on $A$, then it induced actions of $G$ on $l^{\infty}(\mathbb{N},A)$ and on $A_{\infty}$, denoted by $\alpha^{\infty}$ and $\alpha_{\infty}$. Since for any $g\in G$,
@@ 2.2. Cuntz comparison
Below we give some well-known facts about Cuntz comparison. (1) is contained in [31, Proposition 2.6] (and also appears in a slightly different form in the earlier [45, Proposition 2.4]). (2) is [43, Lemma 1.7]. (3) is straightforward (and is also [31, Lemma 2.5(i)]). (4) is [43, Corollary 1.6]. (5) is [31, Lemma 2.8(ii)]. (6) is [31, Lemma 2.8(iii)]. (7) is [31, Lemma 2.9]. (8) is [32, Lemma 2.4(2)]. For (9), it is straightforward to verify that $a$ and $f(a)$ generate the same hereditary subalgebra in $A$. (10) is contained in the discussion following [31, Definition 2.3].
@@
[43, Lemma 2.4] Let $A$ be a simple C*-algebra which is not of type I. Let $a\in A$ be a nonzero positive element, and $l$ a positive integer. Then there exist $l$ mutually orthogonal nonzero positive elements $b_{1},\ldots,b_{l}\in A$ such that $b_{1}\sim_{A}\cdots\sim_{A}b_{l}$, and $b_{1}+\cdots+b_{l}\in\overline{aAa}$.
@@
[43, Lemma 1.5] Let $A$ be a C*-algebra, let $a,b\in A_{+}$, and let $\alpha,\beta\geq 0$. Then
@@
The following lemma is known as Kirchberg’s Slice Lemma, see [44, Lemma 4.1.9].
@@
$\tau$ is called an $n$-quasitrace ($n\geq 2$) if it can be extended to a quasitrace on $M_{n}(A)$. $\tau$ is called normalized if $\|\tau\|=1$.
@@
By replacing normalised 2-quasitraces with densely defined 2-quasitraces, the strict comparison property for positive elements can also be defined for non-unital C*-algebras (see [35]). Recall that an ordered semigroup $S$ is called almost unperforated if for any $x,y\in S$ and $n\in\mathbb{N}$ satisfying $(n+1)x\leq ny$, we have $x\leq y$. By [49, Lemma 6.1], for a simple C*-algebra, having strict comparison of positive elements is equivalent to its Cuntz semigroup being almost unperforated.
@@
The above definition is equivalent to defining strict comparison and the radius of comparison using $\mathrm{W}(A)$; using $\mathrm{Cu}(A)$ yields the same result (see [43, Proposition 6.12]). If $A$ is simple and has finite radius of comparison, then the above infimum is attained, meaning $A$ has $\mathrm{rc}(A)$-comparison (see [50, Proposition 6.3] and [43, Lemma 1.23]).
@@ 2.3. Completely positive order zero map
[55, Definition 2.3] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive map. We say $\varphi$ has order zero if for every $a,b\in A$, we have $\varphi(a)\perp\varphi(b)$ whenever $a\perp b$.
@@
[23, Theorem 2.2, Corollary 4.8] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive map. Then $\varphi$ has order zero if and only if for any $a,b,c\in A$, we have
@@
[55, Theorem 3.3] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive order zero map. Let $C\colon C^{*}(\varphi(A))\subset B$.
Then there exists a positive element $h\in M(C)\cap C^{\prime}$ satisfying $\|h\|=\|\varphi\|$ and a *-homomorphism
@@
[55, Corollary 4.1] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive contractive order zero map. Then, the map defined by $\rho_{\varphi}(\mathrm{id}_{(0,1]}\otimes a)\colon=\varphi(a)$ induces a *-homomorphism $\rho_{\varphi}\colon C_{0}((0,1])\otimes A\to B$.
@@
[55, Corollary 4.2] Let $\varphi\colon A\to B$ be a completely positive contractive order zero map, and let $f\in C_{0}((0,1])$ be a positive function. Let $C$, $h$, and $\pi_{\varphi}$ be as in Theorem 2.16. Then, the map
@@
[55, Corollary 4.3] Let $A,B,C$ and $D$ be C*-algebras and let $\varphi\colon A\to B$ and $\psi\colon C\to D$ be completely positive contractive order zero maps.
Then the induced completely positive contractive map
@@
[55, Corollary 4.4] Let $A$ and $B$ be C*-algebras, $\varphi\colon A\to B$ a completely positive order zero map, and $\tau$ a trace on $B$.
Then, the composition $\tau\circ\varphi$ is a trace.
@@
[55, Corollary 4.5] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive order zero map.
@@ 2.4. Maps under group actions
[29, Definition 1.3] Let $G$ be a compact group, and let $A$, $B$ be C*-algebras. Let $\alpha\colon G\to\mathrm{Aut}(A)$ and $\gamma\colon G\to\mathrm{Aut}(B)$ be actions of $G$ on $A$ and $B$, respectively. Let $F\subseteq A$, $S\subseteq B$, and $\varepsilon>0$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(F,S,\varepsilon)$-approximately equivariant central multiplicative map if:
@@
[37, Definition 1.4] Let $A$ and $B$ be C*-algebras, and let $F\subseteq A$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(n,F,\varepsilon)$-approximately multiplicative map if for any $m\in\{1,2,\ldots,n\}$ and any $x_{1},x_{2},\ldots,x_{m}\in F$, we always have
@@
Let $A$ be a unital C*-algebra, $B$ a C*-algebra, and $\varphi\colon A\to B$ a completely positive map. Let $\varepsilon>0$,
@@ 3. The weak tracial Rokhlin property with comparison
[7, Definition 3.2] Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that $\alpha$ has the weak tracial Rokhlin property if for every finite set $F\subseteq A$, every $\varepsilon>0$, and every $x\in A_{+}$ with $\|x\|=1$, there exist mutually orthogonal contractive positive elements $(d_{g})_{g\in G}\in A$ with $d=\sum_{g\in G}d_{g}$ such that
@@
[6, Lemma 2.14] Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property if and only if for every finite set $F\subseteq A$, every $\varepsilon>0$, and every $x\in A_{+}$ with $\|x\|=1$, there exists a positive element $d\in A^{\alpha}$ with $\|d\|=1$ and contractive positive elements $(d_{g})_{g\in G}\in A$ such that
@@
[37, Definition 2.4] Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the tracial Rokhlin property with comparison if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a projection $p\in A^{\alpha}$ and a unital completely positive map $\psi\colon C(G)\to pAp$ such that
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the weak tracial Rokhlin property with comparison if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
The main technical result of this chapter is Theorem 3.16. This theorem, for given tolerance $\varepsilon>0$ and given compact subsets of $A$ and $A^{\alpha}$, constructs an $\alpha$-invariant contractive positive element $d$ (which is “sufficiently large” in the sense described by conditions (3), (4), (5) of Definition 3.4), and simultaneously constructs a completely positive contractive approximate order zero map from $A$ into $\overline{dA^{\alpha}d}$.
@@
Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every finite set $F\subseteq A$, every $\varepsilon>0$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in A^{\alpha}$ and contractive positive elements $(d_{g})_{g\in G}\in A$ such that
@@
Applying Definition 3.4 with the given $\varepsilon,F,S,x$, and $y$, we obtain a contractive positive element $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every $\varepsilon>0$, every compact set $F\subseteq A$, every compact set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in A^{\alpha}$ and a contractive completely positive equivariant map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
The proof is essentially the same as [37, Lemma 2.9], so we omit it. ∎
The following lemma shows that the element $x$ in condition (6) of Lemma 3.6 can be replaced by another contractive positive element. This will be used in subsequent proofs, so we state it here.
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$, every $z\in A_{+}$ with $\|z\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in A^{\alpha}$ and a contractive completely positive equivariant map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
When $A$ is finite, condition (6) of Definition 3.4, condition (6) of Lemma 3.6, and condition (6) of Lemma 3.7 are redundant. To prove this, we need a technique from [43, Lemma 4.5].
@@
[7, Lemma 2.6] Let $A$ be a unital C*-algebra, let $a,g\in A_{+}$ with $0\leq a,g\leq 1$, and let $\varepsilon_{1},\varepsilon_{2}\geq 0$. Then
@@
The proof is essentially the same as [43, Lemma 2.9]; we simply replace [43, Lemma 1.8] with Lemma 3.8. ∎
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then the action $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
The proof is essentially the same as [43, Lemma 4.5]; we simply replace [43, Lemma 2.9] with Lemma 3.9. ∎
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple separable infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every $x\in A_{+}$ with $\|x\|=1$ and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in(A_{\infty,\alpha}\cap A^{\prime})^{\alpha_{\infty}}$ and an equivariant completely positive contractive order zero map
@@
such that

### v3

@@ Abstract.
In this paper, we introduce compact group actions with the weak tracial Rokhlin property. This concept generalizes both the finite group actions with the weak tracial Rokhlin property and the compact group actions with the tracial Rokhlin property on classifiable C*-algebras (in the sense of the Elliott program). Under this framework, we prove that simplicity, pure infiniteness, tracial $\mathcal{Z}$-stability and the combination of nuclearity and $\mathcal{Z}$-stability can be transferred from the original algebra to the crossed product. We also show that the radius of comparison of the fixed point algebra does not exceed that of the original algebra. Furthermore, we discuss the relationship between our definition and natural generalization of the finite group case in non-Elliott program settings. Finally, we provide a nontrivial example of a compact group action with the weak tracial Rokhlin property with comparison: an action of $(S_{2})^{\mathbb{N}}$ on the Jiang-Su algebra $\mathcal{Z}$. Since $\mathcal{Z}$ contains no nontrivial projections, this action does not possess the tracial Rokhlin property.
@@ 1. Introduction
The Rokhlin property for the case of a single automorphism was originally introduced for von Neumann algebras by Connes in [9]. Later, the Rokhlin property for finite group actions on C*-algebras first appeared in the work of Herman and Jones in [25] and [26]. This property is useful to understand the structure of the crossed product of C*-algebras and properties passing from the original algebra to the crossed product [38]. However, the finite group acitions with the Rokhlin property are rare. Phillips, in [40], introduced the tracial Rokhlin property for finite group actions on unital simple C*-algebras. The tracial Rokhlin property is generic in many cases, and also can be used to study properties passing from the original algebra to the crossed product. Weak versions of the tracial Rokhlin property in which one uses orthogonal positive contractions instead of orthogonal projections were studied for actions on unital simple C*-algebras with few projections [16, 27, 35, 41, 51, 52] (see Definition 3.1). As an example, the flip action on the Jiang-Su algebra $\mathcal{Z}\cong\mathcal{Z}\otimes\mathcal{Z}$ has the weak tracial Rokhlin property but it does not have the tracial Rokhlin property [27]. For the non-unital case, Santiago and Gardella studied the Rokhlin property for finite group actions on non-unital simple C*-algebras in [46] and [18]. Forough and Golestani studied the (weak) tracial Rokhlin property for finite group actions on non-unital simple C*-algebras in [13].
@@ 2.1. Central sequence algebras
let $G$ be a topological group and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action of $G$ on $A$, then it induced actions of $G$ on $l^{\infty}(\mathbb{N},A)$ and on $A_{\infty}$, denoted by $\alpha^{\infty}$ and $\alpha_{\infty}$. Since for any $g\in G$,
@@ 2.2. Cuntz comparison
Below we give some well-known facts about Cuntz comparison. (1) is contained in [30, Proposition 2.6] (and also appears in a slightly different form in the earlier [44, Proposition 2.4]). (2) is [42, Lemma 1.7]. (3) is straightforward (and is also [30, Lemma 2.5(i)]). (4) is [42, Corollary 1.6]. (5) is [30, Lemma 2.8(ii)]. (6) is [30, Lemma 2.8(iii)]. (7) is [30, Lemma 2.9]. (8) is [31, Lemma 2.4(2)]. For (9), it is straightforward to verify that $a$ and $f(a)$ generate the same hereditary subalgebra in $A$. (10) is contained in the discussion following [30, Definition 2.3].
@@
[42, Lemma 2.4] Let $A$ be a simple C*-algebra which is not of type I. Let $a\in A$ be a nonzero positive element, and $l$ a positive integer. Then there exist $l$ mutually orthogonal nonzero positive elements $b_{1},\ldots,b_{l}\in A$ such that $b_{1}\sim_{A}\cdots\sim_{A}b_{l}$, and $b_{1}+\cdots+b_{l}\in\overline{aAa}$.
@@
[42, Lemma 1.5] Let $A$ be a C*-algebra, let $a,b\in A_{+}$, and let $\alpha,\beta\geq 0$. Then
@@
The following lemma is known as Kirchberg’s Slice Lemma, see [43, Lemma 4.1.9].
@@
Definition 2.9.
Let $A$ be a C*-algebra. A linear functional $\tau\colon A\to\mathbb{C}$ is called a trace on $A$ if $\tau$ is positive and satisfies $\tau(xy)=\tau(yx)$ for all $x,y\in A$. If furthermore $\|\tau\|=1$, then $\tau$ is called a tracial state. The set of all tracial states of $A$ is denoted by $\mathrm{T}(A)$.
In fact, for any approximate unit $(e_{\lambda})_{\lambda\in\Lambda}$ of $A$, $\|\tau\|=\lim_{\lambda}\tau(e_{\lambda})$. In particular, if $A$ has a unit, then $\|\tau\|=\tau(1_{A})$.
@@
$\tau$ is called a $n$-quasitrace ($n\geq 2$) if it can be extended to a quasitrace on $M_{n}(A)$. $\tau$ is called normalized if $\|\tau\|=1$.
@@
By replacing normalised 2-quasitraces with densely defined 2-quasitraces, the strict comparison property for positive elements can also be defined for non-unital C*-algebras (see [34]). Recall that an ordered semigroup $S$ is called almost unperforated if for any $x,y\in S$ and $n\in\mathbb{N}$ satisfying $(n+1)x\leq ny$, we have $x\leq y$. By [48, Lemma 6.1], for a simple C*-algebra, having strict comparison of positive elements is equivalent to its Cuntz semigroup being almost unperforated.
@@
The above definition is equivalent to defining strict comparison and the radius of comparison using $\mathrm{W}(A)$; using $\mathrm{Cu}(A)$ yields the same result (see [42, Proposition 6.12]). If $A$ is simple and has finite radius of comparison, then the above infimum is attained, meaning $A$ has $\mathrm{rc}(A)$-comparison (see [50, Proposition 6.3] and [42, Lemma 1.23]).
@@ 2.3. Completely positive order zero map
[53, Definition 2.3] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive map. We say $\varphi$ has order zero if for every $a,b\in A$, we have $\varphi(a)\perp\varphi(b)$ whenever $a\perp b$.
@@
[19, Theorem 2.2, Corollary 4.8] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive map. Then $\varphi$ has order zero if and only if for any $a,b,c\in A$, we have
@@
[53, Theorem 3.3] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive order zero map. Let $C\colon=C^{*}(\varphi(A))\subset B$. Then there exists a positive element $h\in M(C)\cap C^{\prime}$ satisfying $\|h\|=\|\varphi\|$ and a *-homomorphism
@@
[53, Corollary 4.1] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive contractive order zero map. Then, the map defined by $\rho_{\varphi}(\mathrm{id}_{(0,1]}\otimes a)\colon=\varphi(a)$ induces a *-homomorphism $\rho_{\varphi}\colon C_{0}((0,1])\otimes A\to B$.
@@
[53, Corollary 4.2] Let $\varphi\colon A\to B$ be a completely positive contractive order zero map, and let $f\in C_{0}((0,1])$ be a positive function. Let $C$, $h$, and $\pi_{\varphi}$ be as in Theorem 2.17. Then, the map
@@
[53, Corollary 4.3] Let $A,B,C$ and $D$ be C*-algebras and let $\varphi\colon A\to B$ and $\psi\colon C\to D$ be completely positive contractive order zero maps. Then the induced completely positive contractive map
@@
[53, Corollary 4.4] Let $A$ and $B$ be C*-algebras, $\varphi\colon A\to B$ a completely positive order zero map, and $\tau$ a trace on $B$. Then, the composition $\tau\circ\varphi$ is a trace.
@@
[53, Corollary 4.5] Let $A$ and $B$ be C*-algebras and let $\varphi\colon A\to B$ be a completely positive order zero map.
@@ 2.4. Maps under group actions
[28, Definition 1.3] Let $G$ be a compact group, and let $A$, $B$ be C*-algebras. Let $\alpha\colon G\to\mathrm{Aut}(A)$ and $\gamma\colon G\to\mathrm{Aut}(B)$ be actions of $G$ on $A$ and $B$, respectively. Let $F\subseteq A$, $S\subseteq B$, and $\varepsilon>0$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(F,S,\varepsilon)$-approximately equivariant central multiplicative map if:
@@
[36, Definition 1.4] Let $A$ and $B$ be C*-algebras, and let $F\subseteq A$. A completely positive contractive map $\varphi\colon A\to B$ is said to be an $(n,F,\varepsilon)$-approximately multiplicative map if for any $m\in\{1,2,\ldots,n\}$ and any $x_{1},x_{2},\ldots,x_{m}\in F$, we always have
@@
Let $A$ be a unital C*-algebra, $B$ a C*-algebra, and $\varphi\colon A\to B$ a completely positive map. Let $\varepsilon>0$.
@@ 3. The weak tracial Rokhlin property with comparison
[6, Definition 3.2] Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that $\alpha$ has the weak tracial Rokhlin property if for every finite set $F\subseteq A$, every $\varepsilon>0$, and every $x\in A_{+}$ with $\|x\|=1$, there exist mutually orthogonal positive contractions $(d_{g})_{g\in G}\in A$ with $d=\sum_{g\in G}d_{g}$ such that the following hold:
@@
[7, Lemma 2.14] Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property if and only if for every finite set $F\subseteq A$, every $\varepsilon>0$, and every $x\in A_{+}$ with $\|x\|=1$, there exists a positive element $d\in A^{\alpha}$ with $\|d\|=1$ and positive contractions $(d_{g})_{g\in G}\in A$ such that the following hold:
@@
[36, Definition 2.4] Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the tracial Rokhlin property with comparison if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a projection $p\in A^{\alpha}$ and a unital completely positive map $\psi\colon C(G)\to pAp$ such that the following hold:
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the weak tracial Rokhlin property with comparison if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
The main technical result of this chapter is Theorem 3.16. This theorem, for given tolerance $\varepsilon>0$ and given compact subsets of $A$ and $A^{\alpha}$, constructs an $\alpha$-invariant positive contraction $d$ (which is “sufficiently large” in the sense described by conditions (3), (4), (5) of Definition 3.4), and simultaneously constructs a completely positive contractive approximate order zero map from $A$ into $\overline{dA^{\alpha}d}$.
@@
Let $G$ be a finite group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every finite set $F\subseteq A$, every $\varepsilon>0$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in A^{\alpha}$ and positive contractions $(d_{g})_{g\in G}\in A$ such that the following hold:
@@
Applying Definition 3.4 with the given $\varepsilon,F,S,x$, and $y$, we obtain a positive contraction $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every $\varepsilon>0$, every compact set $F\subseteq A$, every compact set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in A^{\alpha}$ and a contractive completely positive equivariant map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
The proof is essentially the same as [36, Lemma 2.9], so we omit it. ∎
The following lemma shows that the element $x$ in condition (6) of Lemma 3.6 can be replaced by another positive contraction. This will be used in subsequent proofs, so we state it here.
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$, every $z\in A_{+}$ with $\|z\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in A^{\alpha}$ and a contractive completely positive equivariant map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
When $A$ is finite, condition (6) of Definition 3.4, condition (6) of Lemma 3.6, and condition (6) of Lemma 3.7 are redundant. To prove this, we need a technique from [42, Lemma 4.5].
@@
[6, Lemma 2.6] Let $A$ be a unital C*-algebra, let $a,g\in A_{+}$ with $0\leq a,g\leq 1$, and let $\varepsilon_{1},\varepsilon_{2}\geq 0$. Then
@@
The proof is essentially the same as [42, Lemma 2.9]; we simply replace [42, Lemma 1.8] with Lemma 3.8. ∎
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then the action $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every $\varepsilon>0$, every finite set $F\subseteq A$, every finite set $S\subseteq C(G)$, every $x\in A_{+}$ with $\|x\|=1$, and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
The proof is essentially the same as [42, Lemma 4.5]; we simply replace [42, Lemma 2.9] with Lemma 3.9. ∎
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple separable infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. Then $\alpha$ has the weak tracial Rokhlin property with comparison if and only if for every $x\in A_{+}$ with $\|x\|=1$ and every $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in(A_{\infty,\alpha}\cap A^{\prime})^{\alpha_{\infty}}$ and an equivariant completely positive contractive order zero map
@@
such that the following hold: